hi im having trouble solving the linear system using substitution can you please help me heres the problem im stuck on y-5=x , 4x-y=4
Answer by ewatrrr(24785)


Hi
solving the linear system using substitution
y-5=x y = x+5
4x-y=4
4x - (x+5) = 4 |substituting for y
solving for x
3x - 5 = 4
3x = 9
x = 3 y = 8 (x+5)
CHECKING our Answer***
y-5 = x
8-5 = 3